# Efficient Low Rank Attention for Long-Context Inference in Large Language Models

### v2

@@ Abstract
As the length of input text grows, the key-value (KV) cache in LLMs imposes prohibitive GPU memory costs and limits long-context inference on resource constrained devices. Existing approaches, such as KV quantization and pruning, reduce memory usage but suffer from numerical precision loss or suboptimal retention of key-value pairs. We introduce Low Rank Query and Key attention (LRQK), a two-stage framework that jointly decomposes the full-precision query and key matrices into compact rank-$r$ factors during the prefill stage, and then uses these low-dimensional projections to compute proxy attention scores in $\mathcal{O}(lr)$ time at each decode step. By selecting only the top-$k$ tokens and a small fixed set of recent tokens, LRQK employs a mixed GPU-CPU cache with a hit-and-miss mechanism that transfers only missing full-precision KV pairs, thereby preserving exact attention outputs while reducing CPU-GPU data movement. Extensive experiments on the RULER and LongBench benchmarks with LLaMA-3-8B and Qwen2.5-7B demonstrate that LRQK matches or surpasses leading sparse-attention methods in long context settings, while delivering significant memory savings with minimal loss in accuracy. Our code is available at https://github.com/tenghuilee/LRQK.
@@ 1 Introduction
Large language models (LLMs) have shown their remarkable capabilities across diverse tasks. Recent LLMs have extended their capabilities to support long-context processing, enabling them to leverage extended sequences for improved performance in applications such as document-level understanding and long-form text generation [1, 2, 3, 4, 5]. Despite these advances, efficient processing of long-contexts presents substantial challenges. The primary constraint stems from computational resource limitations, particularly regarding memory consumption. The KV cache stores historical key–value pairs to prevent redundant computations. However, its size grows linearly with sequence length and quickly imposes prohibitive memory requirements when handling long inputs. Consequently, as context windows expand, the KV cache correspondingly increases in size, eventually becoming the system bottleneck.
@@
While these methods have made significant strides in addressing the KV cache memory bottleneck, they each come with inherent limitations. Quantization methods will sacrifice numerical precision, and the evincing techniques risk eliminating potentially crucial key-value pairs, particularly problematic when tokens deemed unimportant in current contexts later become critical for subsequent reasoning. Pruning methods may inadvertently discard important key-value pairs, leading to suboptimal performance when previously unimportant tokens become critical in subsequent reasoning. Offloading strategies introduce substantial latency overhead due to frequent PCIe data transfers, severely impacting real-time processing capabilities.
@@
To mitigate these challenges, a hybrid attention mechanism is proposed to exploit the intrinsic sparsity of transformer attention patterns while preserving exact computation of attention outputs. At its core lies the observation that full query–key interaction matrix in decoder-only LLMs admits a close low rank approximation. Therefore, the query matrix and key matrix are jointly factorized into compact bases. In the decoding phase, a “proxy” attention score is computed via these factorized query and key matrices, and identify the top-$k$ relevant tokens in the GPU cache. Thus, the ground truth attention score can be carried out on the retrieved subset of $\mathbf{k}$ and $\mathbf{v}$, guaranteeing numerical fidelity. An overview of the proposed method is shown in Figure 1 and the time comparison is represented in Figure 2. The main contributions of this paper can be summarized as:
@@
  • Mixed Cache Management: We implement a sophisticated hybrid GPU-CPU storage system featuring active token retention mechanisms and a hit/miss buffer architecture that minimizes cross-device data transfer. Additionally, a specialized buffer maintains recently accessed keys and values, which empirical analysis confirms consistently receive high attention scores, further optimizing cache hit rates under operational conditions.
The paper is organized as follows. In Section 3, we present two key insights: the low rank nature of the key matrix and the high attention scores between a current token and its neighbors. Section 4 introduces the Low Rank Query and Key (LRQK) method, detailing both prefill and decode stages. In Section 5, we assess the performance of the LRQK method across various models and datasets.
@@ 2.1 Token Eviction
Token eviction methods seek to bound the memory footprint of the KV cache by discarding less informative key–value pairs while preserving generation quality [11, 17, 10, 12, 18, 19]. Early work such as StreamingLLM [11], maintains a fixed “attention sink” of initial tokens together with a sliding window of recent tokens, stabilizing decoding for arbitrarily long sequences. Building on this, H2O [10] formulates eviction as a dynamic submodular optimization problem to retain both recent tokens and “heavy hitters”. Rather than relying on global scoring, SnapKV [12] leverages per-head clustering of attention patterns within a small observation window to select and compress the most relevant KV positions in one pass. More recently, Ada-KV [18] derives a theoretical upper bound on post-eviction attention error and proposes a head-wise adaptive budget allocation strategy, yielding consistent improvements. However, these existing eviction strategies universally trade off either long-range context fidelity or computational simplicity by relying on rigid windows, coarse heuristics, or expensive per-head clustering, therefore resulting in non-negligible approximation error or management overhead. Our approach adopts a joint low‐rank proxy to select only the top-$k$ full-precision KV pairs and a small hit/miss recency cache to load them, guaranteeing exact attention outputs and less CPU-GPU data movement.
@@ 2.2 Dynamic Sparse Attention
Dynamic sparse attention mechanisms represent a significant advancement in transformer-based models by selectively computing attention scores on a subset of tokens while maintaining complete key-value cache storage, thereby substantially reducing computational requirements. Contemporary approaches implement diverse token selection strategies with varying degrees of efficacy: QUEST [20] employs a page-based selection methodology, while Loki [21] leverages low-dimensional attention approximation through principal component analysis (PCA) on key matrices. PALU [22] and LPA [23] also utility the low rank capability in attention. Notably, two recent techniques, InfiniGen [14] and ShadowKV [16] utilize singular value decomposition (SVD) for key-value selection. InfiniGen pre-fetches essential entries using predefined projections via SVD, whereas ShadowKV offloads the entire $\mathbf{V}$ cache to CPU memory after SVD decomposition of the key matrix $(\mathbf{K})$. While effective, these SVD-based approaches incur substantial computational overhead and fundamentally alter the original query/key representations, potentially compromising model fidelity. In contrast, the proposed method combines a joint low rank $\mathbf{Q},\mathbf{K}$ projection and retrieves the relevant full-precision KV pairs on demand, ensuring exact attention computation with less approximation error.
@@ 3.1 Low Rankness of Query and Key
As demonstrated in prior works [16, 14], the key matrix in decoder-only transformers exhibits a pronounced low rank structure. Motivated by this observation, we jointly factorize the full-precision query and key matrices into rank-$r$ components, enforcing
@@ 3.2 Attention Scores Near Current Token
We evaluated the average per‐head scores on the Wikitext-2-v1 test set using both Qwen2.5-7B and LLaMA-3-8B-1M. Results are shown in Figure 4. In every case, the curves exhibit pronounced higher attention scores in the current token and its near neighbors, confirming a strong recency bias. This empirical finding is also observed by StreamingLLM [11], which motivates our inclusion of a compact recency buffer in the GPU cache that persistently holds the most recent tokens, thereby maximizing cache hit rates with minimal overhead.
@@ 4 Proposed Method
The inference process of LLM can be divided into two distinct phases: prefill and decode. When the user provides a prompt, the LLM will first prefill the long prompt once, and then decode the token in an auto-regressive manner. These two phases require specific cache mechanisms, thus we will introduce LRQK for prefill in Section 4.2 and decode in Section 4.3, respectively.
We begin by clarifying the preliminaries and formulation used in our method. In practical implementations, transformer models typically process input tensors with a shape of (batch size, num heads, sequence length, hidden size). However, when methods such as GQA [26] are employed, the number of attention heads used for the query tensor is usually different from those used for the key and value tensors. Therefore, we use the repeated key/value (repeat KV) mechanism, ensuring that the number of heads for the query and the key/value tensors are aligned. In addition, to simplify notation, we also omit the ‘batch size’ and ‘num heads’ dimensions, as computations across different batches and heads are independent. Thus, the query, key, and value matrices as $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{l\times d}$, where $l$ denotes the sequence length and $d$ the hidden size of each head. Moreover, since the prefill and decode process is identical across all transformer layers, we omit the layer indices in our method for clarity.
@@ 4.1 Mixed Caching
To address memory constraints and optimize runtime efficiency, we adopt a mixed caching strategy that leverages both a fast yet limited GPU KV cache and a larger but slower CPU cache. To further reduce the data transfer overhead between GPU and CPU, we introduce a hit-and-miss cache mechanism. In this scheme, if the required tokens are already present in the GPU cache (i.e., a cache hit), they are accessed directly without incurring additional transfer cost. If the required cache is missing in the GPU, only the missing tokens are fetched from the CPU cache, thus avoiding redundant data movement and improving overall system efficiency.
Recognizing that recent tokens are more likely to receive high attention scores (as shown in Figure 4), we further incorporate a dedicated buffer to store the most recent tokens. The GPU KV cache is structured as $\begin{bmatrix}&\text{active tokens}&\text{lite tokens}&\end{bmatrix}$, where the ‘active tokens’ correspond to the top-$k$ tokens corresponding to the set $\Omega_{k}$ with the approximated highest attention scores selected by our low rank approximations of query and key matrices, and the ‘lite tokens’ corresponding to the set $\Omega_{l}$ capture a fixed number of neighboring recent tokens. The total window size of the GPU KV cache is thus the sum of the active and lite tokens, allowing for a balance between relevance and recency in cache composition.
@@ 4.2 Prefill
At this stage, the input consists of a long prompt represented as a matrix $\mathbf{X}\in\mathbb{R}^{l\times d}$, where $l$ denotes the sequence length and $d$ the hidden dimension. By applying linear projections to $\mathbf{X}$, we obtain the query, key, and value matrices: $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{l\times d}$.
@@
where $\mathbf{A}_{Q},\mathbf{A}_{K}\in\mathbb{R}^{l\times r}$ and $r\ll d$ denotes matrix rank. In addition to the joint low rank approximation, we also approximate query and key matrices for memory efficiency:
@@
where $\mathbf{B}_{Q},\mathbf{B}_{K}\in\mathbb{R}^{r\times d}$. Note that the approximation for attention score $\mathbf{QK}^{\top}$ and $\mathbf{Q},\mathbf{K}$ shares the same column/row space spanned by $\mathbf{A}_{Q}$ and $\mathbf{A}_{K}$. For solving (2) and (3), we can formulate it as
@@
To solve the constraint optimization problem (4), we relax it by deriving a Lagrangian:
@@
By solving the Lagrangian (5), with setting the partial derivative to be zero, we can get the solution of the query and key matrices (see Appendix A.1 for details):
@@
From (6), it is notice that the operation of matrix inverse is required. The matrix inverse operation is expensive $O(r^{3})$. However, since the rank $r$ is a small number, and $\mathbf{A}_{K}^{\top}\mathbf{A}_{K},\mathbf{A}_{Q}^{\top}\mathbf{A}_{Q}$, $\mathbf{B}_{K}\mathbf{B}_{K}^{\top},\mathbf{B}_{Q}\mathbf{B}_{Q}^{\top}$ are all with shape $(r,r)$, which are small matrices. Therefore, the computation costs of matrix inverse are relatively low. In addition, when solving $\mathbf{A}_{Q}$ and $\mathbf{A}_{K}$, the memory cost can further be reduced by changing the multiplication order. For example, by changing the order of operations from $(\mathbf{Q}\mathbf{K}^{\top})\mathbf{A}_{K}$ to $\mathbf{Q}(\mathbf{K}^{\top}\mathbf{A}_{K})$, we reduce the computational complexity from $\mathcal{O}(l^{2}d)$ to $\mathcal{O}(rld)$.
@@ 4.3 Decode
Following the prefilling phase, the language model generates the remaining tokens sequentially, adhering to an autoregressive decoding process where each token is predicted based on the previously generated context. Suppose the current stage is $t$, and the input is a row vector $\mathbf{x}_{t}\in\mathbb{R}^{1\times d}$ and its projections are $\mathbf{q}_{t}\in\mathbb{R}^{1\times d}$ and $\mathbf{k}_{t}\in\mathbb{R}^{1\times d}$. In the decode stage, the goal is to compress the $\mathbf{q}_{t}$ and $\mathbf{k}_{t}$ to much smaller ones $\widehat{\mathbf{q}}_{t}\in\mathbb{R}^{1\times r}$ and $\widehat{\mathbf{k}}_{t}\in\mathbb{R}^{1\times r}$, thus the memory footprint of KV cache can be greatly mitigated.
From the previous stage, we obtain the low rank approximations of the query and key matrices in the form of left and right factors: $\mathbf{A}_{Q},\mathbf{A}_{K}\in\mathbb{R}^{l\times r}$, and $\mathbf{B}_{Q},\mathbf{B}_{K}\in\mathbb{R}^{r\times d}$. To maintain consistency in the modeling setup for new tokens, we formulate the following optimization problem for estimating the approximated query and key vectors, $\widehat{\mathbf{q}}_{t}$ and $\widehat{\mathbf{k}}_{t}$, respectively,
@@
The Lagrangian of the optimization problem (7) is written as:
@@
By solving the Lagrangian (8), we obtain the update rules for $\widehat{\mathbf{q}}_{t}$ and $\widehat{\mathbf{k}}_{t}$:
@@
Algorithm 2 summarizes the update rules for $\widehat{\mathbf{q}}_{t}$, $\widehat{\mathbf{k}}_{t}$, $\mathbf{B}_{Q,t}$, and $\mathbf{B}_{K,t}$, the computation of proxy attention, and the update for selected key and value. The graphical representation of the thorough decoding process of LRQK is shown in Figure 1. To avoid random initialization, we use the initial guess of $\widehat{\mathbf{k}}_{t}$ by setting $\lambda_{d1}=0$. The algorithm iterates until convergence or reaches the maximum number of iterations. The attention is computed with selected key and value, $\text{Attention}\left(\mathbf{q}_{t},\mathbf{K}_{\Omega,t},\mathbf{V}_{\Omega,t}\right)$.

### v3

@@ Abstract
As the length of input text increases, the key-value (KV) cache in LLMs imposes prohibitive GPU memory costs and limits long-context inference on resource constrained devices. Existing approaches, such as KV quantization and pruning, reduce memory usage but suffer from numerical precision loss or suboptimal retention of key-value pairs. In this work, Low Rank Query and Key attention (LRQK) is introduced, a two-stage framework that jointly decomposes full-precision query and key matrices into compact rank-$r$ factors during the prefill stage, and then employs these low-dimensional projections to compute proxy attention scores in $\mathcal{O}(lr)$ time at each decode step. By selecting only the top-$k$ tokens and a small fixed set of recent tokens, LRQK employs a mixed GPU-CPU cache with a hit-and-miss mechanism where only missing full-precision KV pairs are transferred, thereby preserving exact attention outputs while reducing CPU-GPU data movement. Extensive experiments on the RULER and LongBench benchmarks with LLaMA-3-8B and Qwen2.5-7B demonstrate that LRQK matches or surpasses leading sparse-attention methods in long context settings, while delivering significant memory savings with minimal accuracy loss. Our code is available at https://github.com/tenghuilee/LRQK.
@@ 1 Introduction
Large language models (LLMs) have shown their remarkable capabilities across diverse tasks. Recent LLMs have extended their capabilities to support long-context processing, enabling them to leverage extended sequences for improved performance in applications such as document level understanding and long-form text generation [1, 2, 3, 4, 5]. Despite these advances, efficient processing of long-contexts presents substantial challenges. The primary constraint stems from computational resource limitations, particularly regarding memory consumption. The KV cache stores historical key-value pairs to prevent redundant computations. However, its size grows linearly with sequence length and quickly imposes prohibitive memory requirements when handling long inputs. Consequently, as context windows expand, the KV cache correspondingly increases in size, eventually becoming the system bottleneck.
@@
While these methods have made significant strides in addressing the KV cache memory bottleneck, they each come with inherent limitations. Quantization methods will sacrifice numerical precision, and the evincing techniques risk eliminating potentially crucial key-value pairs, particularly problematic when tokens deemed unimportant in current contexts later become critical for subsequent reasoning. Pruning methods may inadvertently discard important key-value pairs, leading to suboptimal performance when previously unimportant tokens become critical in subsequent reasoning. Offloading strategies introduce substantial latency overhead due to frequent PCIe data transfers, severely impacting real time processing capabilities.
@@
To mitigate these challenges, a hybrid attention mechanism is proposed to exploit the intrinsic sparsity of transformer attention patterns while preserving exact computation of attention outputs. At its core lies the observation that full query-key interaction matrix in decoder-only LLMs admits a close low rank approximation. Therefore, the query matrix and key matrix are jointly factorized into compact bases. In the decoding phase, a “proxy” attention score is computed via these factorized query and key matrices, and identifies the top-$k$ relevant tokens in the GPU cache. Thus, the ground truth attention score can be carried out on the retrieved subset of $\mathbf{k}$ and $\mathbf{v}$, guaranteeing numerical fidelity. An overview of the proposed method is shown in Figure 1 and the time comparison is represented in Figure 2. The main contributions of this paper can be summarized as:
@@
  • Mixed Cache Management: A sophisticated hybrid GPU-CPU storage system featuring active token retention mechanisms and a hit/miss buffer architecture that minimizes cross device data transfer is implemented. Additionally, a specialized buffer maintains recently accessed keys and values, which empirical analysis confirms consistently receive high attention scores, further optimizing cache hit rates under operational conditions.
The paper is organized as follows. In Section 3, two key insights are presented: the low rank nature of the key matrix and the high attention scores between a current token and its neighbors. Section 4 introduces the Low Rank Query and Key (LRQK) method, detailing both prefill and decode stages. In Section 5, the performance of the LRQK method is assessed across various models and datasets.
@@ 2.1 Token Eviction
Token eviction methods seek to bound the memory footprint of the KV cache by discarding less informative key-value pairs while preserving generation quality [11, 17, 10, 12, 18, 19]. Early work such as StreamingLLM [11], maintains a fixed “attention sink” of initial tokens together with a sliding window of recent tokens, stabilizing decoding for arbitrarily long sequences. Building on this, H2O [10] formulates eviction as a dynamic submodule optimization problem to retain both recent tokens and “heavy hitters”. Rather than relying on global scoring, SnapKV [12] leverages per-head clustering of attention patterns within a small observation window to select and compress the most relevant KV positions in one pass. More recently, Ada-KV [18] derives a theoretical upper bound on post-eviction attention error and proposes a head-wise adaptive budget allocation strategy, yielding consistent improvements. However, these existing eviction strategies universally trade off either long-range context fidelity or computational simplicity by relying on rigid windows, coarse heuristics, or expensive per-head clustering, therefore resulting in non-negligible approximation error or management overhead. Our approach adopts a joint low‐rank proxy to select only the top-$k$ full-precision KV pairs and a small hit/miss recency cache to load them, guaranteeing exact attention outputs and less CPU-GPU data movement.
@@ 2.2 Dynamic Sparse Attention
Dynamic sparse attention mechanisms represent a significant advancement in transformer based models by selectively computing attention scores on a subset of tokens while maintaining complete key-value cache storage, thereby substantially reducing computational requirements. Contemporary approaches implement diverse token selection strategies with varying degrees of efficacy: QUEST [20] employs a page based selection methodology, while Loki [21] leverages low-dimensional attention approximation through principal component analysis (PCA) on key matrices. PALU [22] and LPA [23] also utility the low rank capability in attention. Notably, two recent techniques, InfiniGen [14] and ShadowKV [16] utilize singular value decomposition (SVD) for key-value selection. InfiniGen pre-fetches essential entries using predefined projections via SVD, whereas ShadowKV offloads the entire $\mathbf{V}$ cache to CPU memory after SVD decomposition of the key matrix $(\mathbf{K})$. While effective, these SVD based approaches incur substantial computational overhead and fundamentally alter the original query/key representations, potentially compromising model fidelity. In contrast, the proposed method combines a joint low rank $\mathbf{Q},\mathbf{K}$ projection and retrieves the relevant full-precision KV pairs on demand, ensuring exact attention computation with less approximation error.
@@ 3.1 Low Rankness of Query and Key
As demonstrated in prior works [16, 14], the key matrix in decoder-only transformers exhibits a pronounced low rank structure. Motivated by this observation, the full-precision query and key matrices are jointly factorized into rank-$r$ components, enforcing
@@ 3.2 Attention Scores Near Current Token
The average per-head scores are evaluated on the Wikitext-2-v1 test set using both Qwen2.5-7B and LLaMA-3-8B-1M. Results are shown in Figure 4. In every case, the curves exhibit pronounced higher attention scores in the current token and its near neighbors, confirming a strong recency bias. This empirical finding is also observed by StreamingLLM [11], which motivates our inclusion of a compact recency buffer in the GPU cache that persistently holds the most recent tokens, thereby maximizing cache hit rates with minimal overhead.
@@ 4 Proposed Method
The inference process of LLM can be divided into two distinct phases: prefill and decode. When a prompt is provided by the user, the long prompt is first prefilled once by the LLM, and then the token is decoded autoregressively. Specific cache mechanisms are required for these two phases, the LRQK will be introduced for prefill in Section 4.2 and decode in Section 4.3, respectively.
The preliminaries and formulation used in the proposed method are clarified below. In practical implementations, transformer models typically process input tensors with a shape of (batch size, num heads, sequence length, hidden size). However, when methods such as GQA [26] are employed, the number of attention heads used for the query tensor is usually different from those used for the key and value tensors. Therefore, repeated key/value (repeat KV) mechanism is utilized, ensuring that the number of heads for the query and the key/value tensors are aligned. In addition, to simplify notation, ‘batch size’ and ‘num heads’ dimensions are also omitted, as computations across different batches and heads are independent. Thus, the query, key, and value matrices as $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{l\times d}$, where $l$ denotes the sequence length and $d$ the hidden size of each head. Moreover, since the prefill and decode process is identical across all transformer layers, the layer indices are omitted for clarity.
@@ 4.1 Mixed Caching
To address memory constraints and optimize runtime efficiency, a mixed caching strategy is adopted that leverages both a fast yet limited GPU KV cache and a larger but slower CPU cache. To further reduce the data transfer overhead between GPU and CPU, a hit-and-miss cache mechanism is introduced. In this scheme, when the required tokens are already present in the GPU cache (i.e., a cache hit), direct access is provided without incurring additional transfer cost. When the required cache is missing in the GPU, only the missing tokens are fetched from the CPU cache, therefore avoiding redundant data movement and improving overall system efficiency.
Based on the observation that recent tokens are more likely to receive high attention scores (as shown in Figure 4), a dedicated buffer is incorporated to store the most recent tokens. The GPU KV cache is structured as $\begin{bmatrix}&\text{active tokens}&\text{lite tokens}&\end{bmatrix}$, where the ‘active tokens’ are the top-$k$ tokens corresponding to the set $\Omega_{k}$ with the approximated highest attention scores selected by the low rank approximations of query and key matrices, and the ‘lite tokens’ corresponding to the set $\Omega_{l}$ capture a fixed number of neighboring recent tokens. The total window size of the GPU KV cache is thus the sum of the active and lite tokens, allowing for a balance between relevance and recency in cache composition.
@@ 4.2 Prefill
At this stage, the input consists of a long prompt represented as a matrix $\mathbf{X}\in\mathbb{R}^{l\times d}$, where $l$ denotes the sequence length and $d$ the hidden dimension. By applying linear projections to $\mathbf{X}$, the query, key, and value matrices: $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{l\times d}$ will be obtained.
@@
where $\mathbf{A}_{Q},\mathbf{A}_{K}\in\mathbb{R}^{l\times r}$ and $r\ll d$ denotes matrix rank. In addition to the joint low rank approximation, query and key matrices are approximated for memory efficiency:
@@
where $\mathbf{B}_{Q},\mathbf{B}_{K}\in\mathbb{R}^{r\times d}$. Note that the approximation for attention score $\mathbf{QK}^{\top}$ and $\mathbf{Q},\mathbf{K}$ shares the same column/row space spanned by $\mathbf{A}_{Q}$ and $\mathbf{A}_{K}$. For solving (2) and (3), it can be formulated as
@@
To solve the constraint optimization problem (4), it is relaxed by deriving a Lagrangian:
@@
By solving the Lagrangian (5), with setting the partial derivative to be zero, the solution of the query and key matrices can be obtained (see Appendix A.1 for details):
@@
From (6), it is notice that the operation of matrix inverse is required. The matrix inverse operation is expensive $O(r^{3})$. However, since the rank $r$ is a small number, and $\mathbf{A}_{K}^{\top}\mathbf{A}_{K},\mathbf{A}_{Q}^{\top}\mathbf{A}_{Q}$, $\mathbf{B}_{K}\mathbf{B}_{K}^{\top},\mathbf{B}_{Q}\mathbf{B}_{Q}^{\top}$ are all with shape $(r,r)$, which are small matrices. Therefore, the computation costs of matrix inverse are relatively low. In addition, when solving $\mathbf{A}_{Q}$ and $\mathbf{A}_{K}$, the memory cost can further be reduced by changing the multiplication order. For example, by changing the order of operations from $(\mathbf{Q}\mathbf{K}^{\top})\mathbf{A}_{K}$ to $\mathbf{Q}(\mathbf{K}^{\top}\mathbf{A}_{K})$, the computational complexity is reduced from $\mathcal{O}(l^{2}d)$ to $\mathcal{O}(rld)$.
@@ 4.3 Decode
Following the prefilling phase, the language model generates the remaining tokens sequentially, adhering to an autoregressive decoding process where each token is predicted based on the previously generated context. Suppose the current stage is $t$, and the input is a row vector $\mathbf{x}_{t}\in\mathbb{R}^{1\times d}$ and its projections are $\mathbf{q}_{t}\in\mathbb{R}^{1\times d}$ and $\mathbf{k}_{t}\in\mathbb{R}^{1\times d}$. In the decode stage, the goal is to compress the $\mathbf{q}_{t}$ and $\mathbf{k}_{t}$ to much smaller ones $\widehat{\mathbf{q}}_{t}\in\mathbb{R}^{1\times r}$ and $\widehat{\mathbf{k}}_{t}\in\mathbb{R}^{1\times r}$, therefore the memory footprint of KV cache can be greatly mitigated.
From the previous stage, the low rank approximations of the query and key matrices are obtained in the form of left and right factors: $\mathbf{A}_{Q},\mathbf{A}_{K}\in\mathbb{R}^{l\times r}$, and $\mathbf{B}_{Q},\mathbf{B}_{K}\in\mathbb{R}^{r\times d}$. To maintain consistency in the modeling setup for new tokens, the following optimization problem is formulated for estimating the approximated query and key vectors, $\widehat{\mathbf{q}}_{t}$ and $\widehat{\mathbf{k}}_{t}$, respectively,
@@
The Lagrangian of (7) can be written as:
@@
By solving the Lagrangian (8), update rules for $\widehat{\mathbf{q}}_{t}$ and $\widehat{\mathbf{k}}_{t}$ can be obtained as:
@@
Algorithm 2 summarizes the update rules for $\widehat{\mathbf{q}}_{t}$, $\widehat{\mathbf{k}}_{t}$, $\mathbf{B}_{Q,t}$, and $\mathbf{B}_{K,t}$, the computation of proxy attention, and the update for selected key and value. The graphical representation of the thorough decoding process of LRQK is shown in Figure 1. To avoid random initialization, the initial guess of $\widehat{\mathbf{k}}_{t}$ is utilized by setting $\lambda_{d1}=0$. The algorithm iterates until convergence or reaches the maximum number of iterations. The attention is computed with selected key and value, $\text{Attention}\left(\mathbf{q}_{t},\mathbf{K}_{\Omega,t},\mathbf{V}_{\Omega,t}\right)$.